OpenStudy (anonymous):

Which equation is the inverse variation relationship for the following quantities? y =-12 when x = 5

OpenStudy (anonymous):

the equation for an inverse variation is y=k/x so plug in the number into the variables and you get -12=k/5 solve for k k=-60 then plug k into the equation for an inverse so the equation is y=-60x

OpenStudy (anonymous):

sorry y=-60/x